OpenStudy (anonymous):

convert to rectangular form r=12sintheta-16costheta

zepdrix (zepdrix):

Recall that when switching between rectangular and polar we use:\[\Large\rm x=r \cos \theta,\qquad\qquad y=r \sin \theta\]

zepdrix (zepdrix):

\[\Large\rm r=12\sin \theta-16\cos \theta\]Multiplying each side by r gives us,\[\Large\rm r^2=12r \sin \theta-16 r \cos \theta\]

OpenStudy (anonymous):

i'm at x^2+y^2=12y-16y

OpenStudy (anonymous):

err 16x

zepdrix (zepdrix):

\[\Large\rm x^2+y^2=12y-16x\]Ok great. Move everything to the left side,\[\Large\rm x^2+16x+y^2-12y=0\]And from here we'll want to complete the square on the x's and y's.

OpenStudy (anonymous):

so it'd be (x+8)^2+(y-6)^2=64+36?

zepdrix (zepdrix):

64 for the x's, 36 for the y's, ? Mmm yes very good. Looks like you've got it figured out! :)

OpenStudy (anonymous):

oh... it's that simple?

zepdrix (zepdrix):

Yah they're not too bad. Sometimes they'll make you do some weird stuff with square roots. But for the most part, if you can remember your identities that let you switch, and completing the square (which you seem to be very good at), then they're not too much trouble.

OpenStudy (anonymous):

ok thanks